Adult tickets for a play cost $18 and child tickets cost $2. If there were 20 people at a performance and the theater collected $88 for ticket sales, how many childern attended the play?
Answer by josmiceli(19441)

C = number of children
A = number of adults
A + C = 20
C = 20 - A
18A + 2C = 88
18A + 2(20-A) = 88
18A + 40 - 2A = 88
16A = 48
A = 3
A + C = 20
C = 20 - 3
C = 17
17 children attended the play
check answer
17 + 3 = 20, so 20 people attended
18A + 2C = 88
18*3 + 2*17 = 88
54 + 34 = 88
88 = 88
OK